OpenStudy (anonymous):

Identify the vertex for the graph of y = −2x2 − 12x + 5. (3, −13) (−3, −13) (−3, 23) (3, 23)

OpenStudy (lalaly):

(-3,23)

OpenStudy (anonymous):

vertex given by \[-\frac{b}{2a}=-\frac{-12}{2\times -2}=-3\]